# Getting it Right: Improving Spatial Consistency in Text-to-Image Models

### v1

@@ Abstract
One of the key shortcomings in current text-to-image (T2I) models is their inability to consistently generate images which faithfully follow the spatial relationships specified in the text prompt. In this paper, we offer a comprehensive investigation of this limitation, while also developing datasets and methods that achieve state-of-the-art performance. First, we find that current vision-language datasets do not represent spatial relationships well enough; to alleviate this bottleneck, we create SPRIGHT, the first spatially-focused, large scale dataset, by re-captioning 6 million images from 4 widely used vision datasets. Through a 3-fold evaluation and analysis pipeline, we find that SPRIGHT largely improves upon existing datasets in capturing spatial relationships. To demonstrate its efficacy, we leverage only $\sim$0.25of SPRIGHT and achieve a 22% improvement in generating spatially accurate images while also improving the FID and CMMD scores. Secondly, we find that training on images containing a large number of objects results in substantial improvements in spatial consistency. Notably, we attain state-of-the-art on T2I-CompBench with a spatial score of 0.2133, by fine-tuning on <500 images. Finally, through a set of controlled experiments and ablations, we document multiple findings that we believe will enhance the understanding of factors that affect spatial consistency in text-to-image models. We publicly release our dataset and model to foster further research in this area.
@@ 1 Introduction
In this paper, we present a holistic approach towards investigating and mitigating this shortcoming through diverse lenses. We develop datasets, efficient and effective training techniques, and explore multiple ablations and analyses to understand the behaviour of T2I models towards prompts that contain spatial relationships.
@@
We further demonstrate the impact of SPRIGHT by benchmarking the trade-offs achieved with long and short spatial captions, as well spatially focused and general captions. Finally, we take the first steps towards discovering layer-wise activation patterns associated with spatial relationships, by examining the representation space of CLIP [50] as a text encoder. To summarize, our contributions are as follows:
  1. 1.
@@
  3. 3.
    Through multiple ablations and analyses, we present our diverse findings related to spatial relationships: the impact of long captions, the trade-off between spatial and general captions, layer-wise activations of the CLIP text encoder, effect of training with negations and improvements over attention maps.
@@ 2 Related Work
Text-to-image generative models. Since the initial release of Stable Diffusion [54] and DALL-E [52], different classes of T2I models have been developed, all optimized to generate highly realistic images corresponding to complex natural prompts. Models such as PixArt-Alpha [6], Imagen [55], and ParaDiffusion [62] move away from the CLIP text encoder, and explore traditional language models such as T5 [51] and LLaMA [59] to process text prompts. unCLIP [53] based models have led to multiple methods [48, 33] that leverage a CLIP-based prior as part of their diffusion pipeline. Recent works have also explored improvements in model compression [35, 27] and inference efficiency [41, 42].
@@
Spatial relationships in T2I models. Benchmarking the failures of T2I models on spatial relationships has been well explored by VISOR [20], T2I-CompBench [23], GenEval [17], and DALL-E Eval [9]. Both training-based and test-time adaptations have been developed to specifically improve upon these benchmarks. Control-GPT [68] finetunes a ControlNet [67] model by generating TikZ code representations with GPT-4 and optimizing over grounding tokens to generate images. SpaText [1], GLIGEN[34], and ReCo [64] are additional training-based methods that introduce additional conditioning in their fine-tuning process to achieve better spatial control for image generation. LLM-Grounded Diffusion [36] is a test-time multi-step method that improves over layout generated LLMs in an iterative manner. Layout Guidance [7] restricts objects to their annotated bounding box locations through refinement of attention maps during inference. LayoutGPT [15] creates initial layout guidance from LLMs in the form of CSS code, and then uses layout-to-image models to create indoor scenes.
@@
Synthetic captions for T2I models. The efficacy of using descriptive and detailed captions has recently been explored by DALL-E 3 [45], PixArt-Alpha [6] and RECAP [58]. DALL-E 3 builds an image captioning module by jointly optimizing over a CLIP and language modeling objective. RECAP fine-tunes an image captioning model (PALI [8]) and reports the advantages of fine-tuning the Stable Diffusion family of models on long, synthetic captions. PixArt-Alpha also re-captions images from the LAION [56] and Segment Anything [28] datasets; however their key focus is to develop descriptive image captions. On the contrary, our goal is to develop captions that explicitly capture the spatial relationships seen in the image.
@@
3 The SPRIGHT Dataset
3.1 Creating the SPRIGHT Dataset
We re-caption $\sim$6M images from 4 well-known VL datasets:
  • CC-12M [4] : We re-caption a total of 2.3 million images from the CC-12M dataset, filtering out images of resolution less than 768.
  • Segment Anything (SA) [28] : We select Segment Anything as we find that most images in it encapsulates a large number of objects; i.e. larger number of spatial relationships can be captured from a given image. We re-caption 3.5 million images as part of our re-captioning process. Since SA does not have ground-truth captions, we also generate its general captions using the CoCa [65] model.
  • COCO [37] : We re-caption images from the COCO Validation Set, totalling to $\sim$ 40,000 images.
  • LAION-Aesthetics [^2] https://laion.ai/blog/laion-aesthetics/ : We consider 50,000 images from the LAION-Aesthetics dataset and do not include more images as the dataset is currently publicly unavailable. (We also do not release the upsampled captions at this time, for the same reason.)
We leverage LLaVA-1.5-13B [38] to produce synthetic spatial captions, and use the following prompt to create the SPRIGHT Dataset:
We also explore other relevant open models such as FUYU-8B [2] and IDEFICS-9B-Instruct [32], but find LLaVA to be the most consistent at generating spatially accurate captions.
@@ 3.2 Impact of SPRIGHT
As shown in Table 1, we find that SPRIGHT greatly enhances the presence of spatial phrases across all relationship types, on all the datasets. Specifically, considering 11 spatial and relational relationships, ground-truth captions of COCO and LAION only capture 21.05% and 6.03% of relationships; on the other hand, SPRIGHT captures 304.79% and 284.7%, respectively. This shows that captions in VL datasets largely lack the presence of spatial relationships, and that SPRIGHT is able to improve upon this shortcoming by almost always capturing spatial relationships in every sentence.
Our captions offer several improvements beyond the spatial aspects: (i) As depicted in Table 2 we improve the overall linguistic quality compared to the original captions, and (ii) we identify more objects and amplify their occurrences as illustrated in Figure 2; where we plot the top 10 objects present in the original COCO Captions and find that we significantly upsample their corresponding presence in SPRIGHT.
@@
Nouns Adjective Verbs Tokens
COCO 3.00 0.83 0.04 11.28
   + SPRIGHT 14.31 3.82 0.15 68.22
CC-12M 3.35 1.36 0.26 22.93
   + SPRIGHT 13.99 4.36 0.16 67.41
LAION 1.78 0.70 0.11 12.49
   + SPRIGHT 14.32 4.53 0.14 69.74
Segment Anything 3.1 0.79 0.01 9.88
   + SPRIGHT 13.42 4.65 0.12 63.9
Table 2: In addition to improving the presence of spatial relationships, SPRIGHT also enhances the overall linguistic variations and diversity of captions in comparison to their original versions.
Figure 2: (Left) On comparing word cloud representations on the COCO dataset, we find that SPRIGHT significantly amplifies the presence of spatial relationships. (Right) Correspondingly, we also improve fine-grained object identification and their presence in our captions.
Figure 3: The distribution of ratings assigned to 100 captions from SPRIGHT by GPT-4(V), on a scale from 1 to 10.
@@ 3.3 Dataset Validation
We perform 3 levels of evaluation to validate the correctness of the captions present in SPRIGHT:
Category # Examples Accuracy (%)
Overall 88.9
FAITHScore
Entities 149,393 91.4
Relations 167,786 85.8
Colors 10,386 83.1
Counting 59,118 94.5
Other 29,661 89.0
Spatial 45,663 83.6
Table 3: FAITHScore caption evaluation of our SPRIGHT dataset. On a sample of 40,000 random captions, SPRIGHT obtains an 88.9% accuracy, comparable with the reported 86% and 94% on LLaVA-1k and MSCOCO-Captions, respectively [25]. On the subset of atomic claims about spatial relations, SPRIGHT is correct 83.6% of the time.
1. FAITHScore. Following [25], we leverage a large language model to deconstruct generated captions into atomic (simple) claims that can be individually and independently verified in a Visual Question Answering (VQA) format. We randomly sample 40,000 image-generated caption pairs from our dataset, and prompt GPT-3.5-Turbo to identify descriptive phrases (as opposed to subjective analysis that cannot be verified from the image) and decompose the descriptions into atomic statements. These atomic statements are then passed to LLaVA-1.5-13B for verification, and correctness is aggregated over 5 categories: entity, relation, colors, counting, and other attributes. We also measure correctness on spatial-related atomic statements, i.e., those containing one of the keywords left/right, above/below, near/far, large/small and background/foreground. As shown in Table 3 the captions are on average 88.9% correct, with spatially-focused relations, being 83.6% correct. Since there is some uncertainty about bias induced by using LLaVA to evaluate LLaVA-generated captions, we also verify the caption quality in other ways, as described next.
2. GPT-4 (V). Inspired by recent methods [45, 71], we perform a small-scale study on 100 images to evaluate our captions with GPT-4(V) [46]. We prompt GPT-4(V) to rate each caption between a score of 1 to 10, especially focusing on the correctness of the spatial relationships captured. We present the distribution of scores in Figure 3 – the mean and median rating is 6.41 and 7.00 respectively. We present the GPT-4 (V) prompt used in Appendix A.
@@ 4 Improving Spatial Consistency
In this section, we leverage SPRIGHT in an effective and efficient manner, and describe methodologies that significantly advance spatial reasoning in T2I models. Our experimental setting is as follows: Our base model is Stable Diffusion v2.1  [^3] https://huggingface.co/stabilityai/stable-diffusion-2-1; our training and validation set consists of 13,500 and 1,500 images respectively, randomly sampled in a 50:50 split between LAION-Aesthetics and Segment Anything. Each image is paired with both, a general and a spatial caption (from SPRIGHT). During fine-tuning, for each image, we randomly choose one of the given caption types in a 50:50 ratio. We fine-tune the U-Net and the CLIP Text Encoder as part of our training, both with a learning rate $5\times 10^{-6}$ optimized by AdamW [40] and a global batch size of 128. While we train the U-Net for 15,000 steps, the CLIP text encoder remains frozen during the first 10,000 steps. We develop our code-base on top of the Diffusers library [60].
@@
Table 4: Quantitative metrics across multiple spatial reasoning and image fidelity metrics, demonstrating the effectiveness of high quality spatially-focused captions in SPRIGHT. Green indicates results of the model fine-tuned on SPRIGHT. To calculate FID, we use cfg = 3.0 and 7.0 for the baseline and the fine-tuned model, respectively
Figure 4: Illustative examples from our model, as described in Section 4.1, on evaluation prompts from the T2I-CompBench benchmark.
@@ 4.1 Improving upon Baseline Methods
We present our quantitative results on the spatial relationship benchmarks (VISOR [20], T2I-CompBench [23]) and image fidelity metrics in Table 4. To account for the inconsistencies associated with FID [10, 47], we also report results on CMMD [24]. Across all metrics, we significantly improve upon the base model by fine-tuning on <15k images. We conclude that the dense, spatially focused captions in SPRIGHT provide effective spatial guidance to T2I models, and alleviate the need to scale up fine-tuning on a large number of images. Illustrative examples on evaluation prompts from T2I-CompBench are shown in Figure 4, while images corresponding to complex prompts are presented in Figure 5. As shown in Figure 5, we find that the model captures complex spatial relationships (top right), relative sizes (large) and patterns (swirling).
@@ 4.2 Efficient Training Methodology
We partition our dataset of <15k images, as described in Section 4, and bucket them into multiple splits, where each split has images constrained with the # of objects present in them. We achieve this by leveraging the open-world image tagging model Recognize Anything [69]. Subsequently, we train respective models using training images solely from a given split, and benchmark their performance. We keep the same hyper-parameters as before, only initiating training of the CLIP Text Encoder from the beginning. For our dataset split of <15k images, we find the median # of objects/ image to be 11. Thus, we create 5 splits, train corresponding models and benchmark them in Table 5. We find that with the increase in the # of objects / image, we are iteratively able to improve upon the spatial fidelity.
[…]
<6 444 0.1309
<11 1346 0.1468
11 1346 0.1667
>11 1346 0.1613
> 18 444 0.2133
@@
Table 6: Comparing baseline SD 2.1 with our state-of-the-art model, across multiple spatial reasoning and image fidelity metrics, as described in Section 4.2 Green indicates results from our model. To calculate FID, we use cfg = 3.0 and 7.5 for baseline and our model, respectively
@@
Table 8: Results on the GenEval Benchmark. In addition to improving spatial relationships, we also alleviate performance while generating 1/2 objects along with generating the correct number of objects.
Our major finding is that, with 444 training images and spatial captions from SPRIGHT, we are able to attain a 41% improvement over baseline SD 2.1 and achieve state-of-the art across all reported models on the T2I-CompBench spatial score. In Table 6, we find that compared to SD 2.1, we largely improve on all aspects of the VISOR score, while improving the ZS-FID and CMMD score on COCO-30K images by 25.39% and 27.16%, respectively. On VISOR, our key findings are : (a) We increase the Object Accuracy (OA) score by 26.86%, indicating that we are much better at generating objects mentioned in the input prompt, (b) Our VISOR4 score of 16.15% denotes that for a given input prompt, we consistently generate a spatially accurate image (Figure 6). Our model outperforms existing methods on the VISOR benchmark (Table 7), with a notable improvement of 115.61% improvement on the VISOR4, denoting superior consistency in generating spatially correct images.
Next, we compare our model’s performance on the GenEval [17] benchmark (Table 8), and find that in addition to improving spatial relationship (see Position), we also enhance the ability to generate 1 and 2 objects, along with generating the correct number of objects. This further emphasizes our model’s ability to correctly generate objects mentioned in the input prompt. We present additional results on T2I-CompBench in Appendix B.
Figure 6: Generated images from our model, as described in Section 4.2, on evaluation prompts from T2I-CompBench. We find that for a given text prompt, our model consistently generates spatially accurate images.
Throughout our experiments, we find that our training approach not only preserves but also enhances the non-spatial aspects associated with a text-to-image model. We present illustrative images on VISOR prompts in Appendix C.
5 Ablations and Analyses
To fully ascertain the impact of spatially-focused captions in SPRIGHT, we experiment with multiple nuances of our dataset and the corresponding T2I pipeline. Unless stated otherwise, the experimental setup is the same as Section 4.
Figure 7: Illustrative comparisons between models trained on varying ratio of spatial experiments. We find that when trained on 50% and 75% ratio of spatial captions, model achieve their optimal performance. For each prompt, the images are sampled from the same seed, for a fair comparison.
5.1 Optimal Ratio of Spatial Captions
To understand the impact of spatially focused captions in comparison to ground-truth (non-spatial) captions, we fine-tune different models by varying the % of spatial captions as (25/50/75/100). Our results suggest that the model trained on 50% spatial captions achieves the most optimal scores on T2I-CompBench (Table 9). We find that models trained on only 25% of spatial captions suffer largely from incorrect spatial relationships whereas the model trained only on spatial captions fails to generate both the mentioned objects in the input prompt. Figure 7 shows illustrative examples.
@@ 5.1 Optimal Ratio of Spatial Captions
Table 9: Comparing the T2I-CompBench Spatial Scores between models trained on varying ratio of spatial captions. Fine-tuning on a ratio of 50% and 75% of spatial captions yield the most optimal results on spatial relationships.
5.2 Impact of Long and Short Spatial Captions
We also compare the effect of fine-tuning with shorter and longer variants of spatial captions. We create the shorter variants by randomly sampling 1 sentence from the longer caption, and fine-tune multiple models, with different setups. Across, all setups, we find (Table 10) that longer captions perform better than their shorter counterparts. In fact, we find that CLIP fine-tuning hurts performance while using shorter captions, but has a positive impact on longer captions. This potentially happens because fine-tuning CLIP enables T2I models to generalize better to longer captions, which are out-of-distribution at the onset of training by virtue of them being initially pre-trained on short(er) captions from datasets such as LAION.
@@ 5.2 Impact of Long and Short Spatial Captions
Table 10: Comparing the T2I-CompBench Spatial Scores between models trained on long and short spatial captions. Across multiple fine-tuning setups, we find that longer spatial captions lead to better improvements in spatial consistency.

### v2

@@ Abstract
One of the key shortcomings in current text-to-image (T2I) models is their inability to consistently generate images which faithfully follow the spatial relationships specified in the text prompt. In this paper, we offer a comprehensive investigation of this limitation, while also developing datasets and methods that support algorithmic solutions to improve spatial reasoning in T2I models. We find that spatial relationships are under-represented in the image descriptions found in current vision-language datasets. To alleviate this data bottleneck, we create SPRIGHT, the first spatially focused, large-scale dataset, by re-captioning 6 million images from 4 widely used vision datasets and through a 3-fold evaluation and analysis pipeline, show that SPRIGHT improves the proportion of spatial relationships in existing datasets. We show the efficacy of SPRIGHT data by showing that using only $\sim$0.25of SPRIGHT results in a 22% improvement in generating spatially accurate images while also improving FID and CMMD scores. We also find that training on images containing a larger number of objects leads to substantial improvements in spatial consistency, including state-of-the-art results on T2I-CompBench with a spatial score of 0.2133, by fine-tuning on <500 images. Through a set of controlled experiments and ablations, we document additional findings that could support future work that seeks to understand factors that affect spatial consistency in text-to-image models. Project page : https://spright-t2i.github.io/
Keywords:
Text to Image Generation Spatial Relationships
@@ 1 Introduction
In this paper, we present a holistic approach towards investigating and mitigating this shortcoming through diverse lenses. We develop datasets, efficient training techniques, and explore multiple ablations and analyses to understand the behaviour of T2I models towards prompts that contain spatial relationships.
@@
We further demonstrate the impact of SPRIGHT by benchmarking the trade-offs achieved with long and short spatial captions, as well as spatially focused and general captions. We take the first steps towards discovering layer-wise activation patterns associated with spatial relationships, by examining the representation space of CLIP [46] as a text encoder.
Our contributions and key findings are summarized below:
@@
  • Through multiple ablations and analyses, we present our findings related to spatial relationships: the impact of long captions, the trade-off between spatial and general captions, layer-wise activations of the CLIP text encoder, effect of training with negations and improvements over attention maps.
@@ 2 Related Work
Since the initial release of Stable Diffusion [50] and DALL-E [49], different classes of T2I models have been developed, all optimized to generate highly realistic images corresponding to complex natural language prompts. Models such as PixArt-Alpha [6], Imagen [51], and ParaDiffusion [56] move away from the CLIP text encoder, and explore traditional language models such as T5 [47] and LLaMA [54] to process text prompts. unCLIP [48] based models have led to multiple methods [43, 30] that leverage a CLIP-based prior as part of their diffusion pipeline.
Spatial relationships in T2I models.
@@
Benchmarking the failures of T2I models on spatial relationships has been well explored by VISOR [20], T2I-CompBench [23], GenEval [17], and DALL-E Eval [9]. Both training-based and test-time adaptations have been developed to specifically improve upon these benchmarks. Control-GPT [62] finetunes a ControlNet [61] model by generating TikZ code representations with GPT-4 and optimizing over grounding tokens to generate images. SpaText [1], GLIGEN[31], and ReCo [58] are training-based methods that introduce additional conditioning in their fine-tuning process to achieve better spatial control for image generation. LLM-Grounded Diffusion [32] is a test-time multi-step method that improves over layout generated LLMs in an iterative manner. Layout Guidance [7] restricts objects to their annotated bounding box locations through refinement of attention maps during inference. LayoutGPT [15] creates an LLM guided initial layout in the form of CSS, and then uses layout-to-image models to create indoor scenes. REVISION [4] leverages 3D rendering engines to generate synthetic images which act as additional guidance during image synthesis for accurate depiction of spatial relationships in the generated image.
Synthetic captions for T2I models.
@@
The efficacy of using descriptive and detailed captions has recently been explored by DALL-E 3 [40], PixArt-Alpha [6] and RECAP [53]. DALL-E 3 builds an image captioning module by jointly optimizing over a CLIP and language modeling objective. RECAP fine-tunes an image captioning model (PALI [8]) and reports the advantages of fine-tuning the Stable Diffusion family of models on long, synthetic captions. PixArt-Alpha also re-captions images from the LAION [52] and Segment Anything [26] datasets; however their key focus is to develop descriptive image captions. On the contrary, our goal is to develop captions that explicitly capture the spatial relationships seen in the image.
3 The SPRIGHT Dataset
3.1 Creating the SPRIGHT Dataset
We re-caption approximately six million images from four existing vision-language datasets, i.e. datasets containing images and their corresponding natural language descriptions:
  • CC-12M [3] : We re-caption a total of 2.3 million images from the CC-12M dataset, filtering out images of resolution less than 768 $\times$ 768.
  • Segment Anything (SA) [26] : We select Segment Anything as most images in it encapsulates a large number of objects; i.e. larger number of spatial relationships can be captured from a given image. We re-caption 3.5 million images as part of our re-captioning process. Since SA does not have ground-truth captions, we generate its general captions using the CoCa [59] model.
  • COCO [33] : We re-caption images ($\sim$ 40,000) from the validation set.
  • LAION-Aesthetics [^2] https://laion.ai/blog/laion-aesthetics/ : We used 50,000 images from LAION-Aesthetics.  [^3] The entire LAION-5B dataset has been recalled for safety review: https://laion.ai/notes/laion-maintenance/. We will release our re-captioning outputs for these images based on the conclusions of this safety review.
We use LLaVA-1.5-13B [34] with the following prompt to produce synthetic spatial captions to create the SPRIGHT dataset:
@@ 3.1 Creating the SPRIGHT Dataset
Figure 2: Compared to ground truth COCO captions,(Left) Word cloud representations showing that SPRIGHT captions significantly amplify the presence of spatial relationships. (Right)SPRIGHT captions also capture a higher number of object occurances.
@@ 3.2 Impact of SPRIGHT
Table 1 shows that SPRIGHT enhances the presence of spatial phrases across all relationship types on all the datasets.
Table 2: In addition to improving the presence of spatial relationships, SPRIGHT enhances linguistic diversity of captions in comparison to their original versions.
@@
Nouns Adjectives Verbs Tokens
COCO $\rightarrow$ COCO+SPRIGHT 3.00 $\rightarrow$ 14.31 0.83 $\rightarrow$ 3.82 0.04 $\rightarrow$ 0.15 11.28 $\rightarrow$ 68.22
CC-12M $\rightarrow$ CC-12M+SPRIGHT 3.35 $\rightarrow$ 13.99 1.36 $\rightarrow$ 4.36 0.26 $\rightarrow$ 0.16 22.93 $\rightarrow$ 67.41
LAION $\rightarrow$ LAION+SPRIGHT 1.78 $\rightarrow$ 14.32 0.70 $\rightarrow$ 4.53 0.11 $\rightarrow$ 0.14 12.49 $\rightarrow$ 69.74
SA $\rightarrow$ SA+SPRIGHT 3.10 $\rightarrow$ 13.42 0.79 $\rightarrow$ 4.65 0.01 $\rightarrow$ 0.12 09.88 $\rightarrow$ 63.90
For 11 relationships, while the ground-truth captions of COCO and LAION only capture 21.05% and 6.03% of relationships, SPRIGHT captures 304.79% and 284.7%, respectively, i.e. each re-captioned COCO image in SPRIGHT has $\sim$3 spatial phrases. This shows that captions in VL datasets largely lack the presence of spatial relationships, and that SPRIGHT is able to improve upon this shortcoming by almost always capturing spatial relationships in every sentence. Our captions offer several improvements beyond the spatial aspects: (i) As depicted in Table 2 we improve the overall linguistic quality compared to the original captions, and (ii) we identify more objects and amplify their occurrences as illustrated in Figure 2; where we plot the top 10 objects present in the original COCO Captions and find that we significantly upsample their corresponding presence in SPRIGHT.
@@ 3.3 Dataset Validation
We perform 3 levels of evaluation to validate the SPRIGHT captions:
1. FAITHScore. Following [25], we leverage a large language model to deconstruct generated captions into atomic (simple) claims that can be individually and independently verified in a Visual Question Answering (VQA) format. We randomly sample 40,000 image-generated caption pairs from our dataset, and prompt GPT-3.5-Turbo to identify descriptive phrases (as opposed to subjective analysis that cannot be verified from the image) and decompose the descriptions into atomic statements. These atomic statements are then passed to LLaVA-1.5-13B for verification, and correctness is aggregated over 5 categories: entity, relation, colors, counting, and other attributes. We also measure correctness on spatial-related atomic statements, i.e., those containing one of the keywords left/right, above/below, near/far, large/small and background/foreground. The captions are on average 88.9% correct, with spatially-focused relations, being 83.6% correct; with the detailed breakdown presented in the Supplementary Materials. Since there is some uncertainty about bias induced by using LLaVA to evaluate LLaVA-generated captions, we also verify the caption quality in other ways, as described next.
2. GPT-4 (V). Inspired by recent methods [40, 65], we perform a small-scale study on a split of 444 images from LAION and SA (from Section 4.2) to evaluate our captions with GPT-4(V) Turbo [41]. We prompt GPT-4(V) to rate each caption between a score of 1 to 10, especially focusing on the correctness of the spatial relationships captured. Captions of images from LAION and SA had a {mean, median} rating of {7.49,8} and {7.36,8}, respectively. We present the prompt used in the Supplementary Materials.
@@ 4 Improving Spatial Consistency
In this section, we leverage SPRIGHT in an effective and efficient manner, and describe methodologies that significantly advance spatial reasoning in T2I models. We use Stable Diffusion v2.1  [^4] https://huggingface.co/stabilityai/stable-diffusion-2-1 as the base model and our training and validation set consists of 13,500 and 1,500 images respectively, randomly sampled in a 50:50 split between LAION-Aesthetics and Segment Anything. Each image is paired with a typical caption and a spatial caption (from SPRIGHT). During fine-tuning, for each image, we randomly choose one of the given caption types in a 50:50 ratio. We fine-tune the U-Net and the CLIP text encoder as part of our training, both with a learning rate $5\times 10^{-6}$ optimized by AdamW [36] and a global batch size of 128. While we train the U-Net for 15,000 steps, the CLIP text encoder remains frozen during the first 10,000 steps. We develop our code-base on top of the Diffusers library [44].
Table 3: Quantitative metrics across multiple spatial reasoning and image fidelity metrics, demonstrating the effectiveness of high quality spatially-focused captions in SPRIGHT. Green indicates results of the model fine-tuned on SPRIGHT. For FID, we use cfg = 3.0 and 7.0 for the baseline and the fine-tuned model, respectively.
@@ 4.1 Improving upon Baseline Methods
We present results on the spatial relationship benchmarks (VISOR [20], T2I-CompBench [23]) and image fidelity metrics in Table 3. To account for the inconsistencies associated with FID [10, 42], we also report results on CMMD [24]. Across all metrics, our method significantly improves upon the base model by fine-tuning on <15k images. We conclude that the dense, spatially focused captions in SPRIGHT provide effective spatial guidance to T2I models, and alleviate the need to scale up fine-tuning on a large number of images. As shown in Figure 3, the model captures complex spatial relationships (top right), relative sizes (large) and patterns (swirling).
# of Objects per Image <6 <11 11 >11 > 18
# of Training Images 444 1346 1346 1346 444
T2I-CompBench Spatial Score ($\uparrow$) 0.1309 0.1468 0.1667 0.1613 0.2133
@@ 4.2 Efficient Training Methodology
For our dataset of <15k images the median # of objects/image = 11. We partition our dataset into multiple subsets based on the maximum number of objects present in an image.
Table 5: Comparing baseline SD 2.1 with our state-of-the-art model, across multiple spatial reasoning and image fidelity metrics, as described in Section 4.2. Green indicates results from our model. For FID, we use cfg = 3.0 and 7.5 for the baseline model and our model, respectively
@@
This partitioning is automated using the open-world image tagging model Recognize Anything [63]. We create five subsets, train corresponding models on a single subset and benchmark them in Table 4. We keep the same hyper-parameters as before, only initiating training of the CLIP Text Encoder from the beginning. With an increase in the # of objects / image, iterative improvement in spatial fidelity is observed, with the best score for the subset containing greater than 18 objects.
Our major finding is that, with 444 training images and spatial captions from SPRIGHT, we achieve a 41% improvement over the baseline SD 2.1 and attain state-of-the-art performance across all reported models on the T2I-CompBench spatial score. In Table 5, compared to SD 2.1, we significantly improve all aspects of the VISOR score, while also enhancing the ZS-FID and CMMD scores on COCO-30K images by 25.39% and 27.16%, respectively. Our key findings on VISOR (Table 6) include: (a) a 26.86% increase in the Object Accuracy (OA) score, indicating substantial gains in generating objects mentioned in the input prompt, and (b) a VISOR4 score of 16.15%, demonstrating our model’s consistent generation of spatially accurate images.
@@
Table 7: Results on the GenEval Benchmark. In addition to spatial relationships, we also improve model performance in generating the correct number of objects.
@@
We also compare our model’s performance on the GenEval [17] benchmark (Table 7), and find that in addition to improving spatial relationship (see Position), our model shows improvement in generating 1 and 2 objects, along with the correct number of objects. Throughout our experiments, our training approach not only preserves but also enhances the non-spatial aspects associated with a text-to-image model. Additional results and illustrations from VISOR and T2I-CompBench are provided in the Supplementary Materials.
Table 8: Comparing (a) the effect the percentage of spatial captions and (b) the effect of long and short spatial captions.
@@
5 Ablation Studies and Analyses
To fully ascertain the impact of spatially-focused captions in SPRIGHT, we experiment with multiple nuances of our dataset and the corresponding T2I pipeline. Unless stated otherwise, the experimental setup identical to Section 4.
Figure 4: Illustrative comparisons between models trained on varying ratio of spatial experiments. Models trained on 50% and 75% spatial captions are optimal.
5.1 Optimal Ratio of Spatial Captions
To understand the impact of spatially focused captions in comparison to ground-truth captions, we fine-tune different models by varying the % of spatial captions. The results suggest that the model trained on 50% spatial captions achieves the best spatial scores on T2I-CompBench (Table 3(b) (a)). The models trained on only 25% of spatial captions suffer largely from incorrect spatial relationships whereas the model trained only on spatial captions fails to generate the mentioned objects in the input prompt. Figure 4 shows illustrative examples.
5.2 Impact of Long and Short Spatial Captions
We also compare the effect of fine-tuning with shorter and longer variants of spatial captions. We create the shorter variants by randomly sampling 1 sentence from the longer caption, and fine-tune multiple models, with different setups. Across, all setups, (Table 3(b) (b)) longer captions perform better than their shorter counterparts. In fact, CLIP fine-tuning hurts performance while using shorter captions, but has a positive impact on longer captions. This potentially happens because fine-tuning CLIP enables T2I models to generalize better to longer captions, which are out-of-distribution at the onset of training as they are initially pre-trained on short(er) captions from datasets such as LAION.
Figure 5: Comparison of layer-wise representations between Baseline CLIP (X-axis) and fine-tuned CLIP on SPRIGHT (Y-axis). Spatial captions show distinct representations in output attention projections and MLP layers, while layer norm layers are more similar. The representation gap widens with long, complex prompts, suggesting spatial prompts in SPRIGHT create diverse embeddings.
Table 9: CLIP fine-tuned on SPRIGHT is able to differentiate the spatial nuances present in a textual prompt. While Baseline CLIP shows a high similarity for spatially different prompts, SPRIGHT enables better fine-grained understanding.
“above” “below” “to the left of” “to the right of” “in front of” “behind”
Baseline CLIP 0.9225 0.9259 0.9229 0.9223 0.9231 0.9289
CLIP + SPRIGHT 0.8674 0.8673 0.8658 0.8528 0.8417 0.8713